OpenStudy (anonymous):

Does a limit that gives X/0 ( x being any number except than 0 ) will be infinity ( negative or positive depending on x) ? Or can it also not exist at all?

OpenStudy (anonymous):

I think it doesn't exist, but I guess you can interpretate as positive infinite to x>0 and negative infinite to x<0

OpenStudy (anonymous):

Or it depends upon your definition of limit?

OpenStudy (anonymous):

I had this problem here lim x--> 2 on the left hand side (3x+4) ^2 / x ^2-x-2 gave me - infinity. Denominator factors out to x+1 & x-2. So it cannot be cancelled out in someway. The answer key gives me a negative. thing is I don't understand why some problems do that :/

OpenStudy (anonymous):

It's -infinite to \[x \rightarrow2^-\] and +infinite to -infinite to \[x \rightarrow2^+\] http://tutorial.math.lamar.edu/Classes/CalcI/InfiniteLimits.aspx

OpenStudy (anonymous):

thanks :) ill read it ill give you a best response :)

OpenStudy (anonymous):

I had a mistake while writting the response. Is -infinite when \[x \rightarrow 2^-\] And +infinite when: \[x \rightarrow 2^+\] Ignore the "-infinite to" after the "+infinite to"

OpenStudy (unklerhaukus):

\[\lim\limits_{x\rightarrow 2}\quad\frac{ (3x+4) ^2 }{ x ^2-x-2}\]\[=\lim\limits_{x\rightarrow 2}\quad\frac{ (3x+4) ^2 }{ (x+1)(x-2)}\]

hartnn (hartnn):

it all depends on values of left limit and right limit. If left limit = right limit = + infinity then limit EXIST and equals +infinity If left limit = right limit = - infinity then limit EXIST and equals -infinity if left limit NOT equal to right limit, then limit DOES NOT EXIST. no matter what the values of left and right limit are. In your case, u had (x-2) in denominator. for left limit x->2- means x is very very near to -2 but less. (like -1.99999) means x-2 will be negative. Hence, the answer to left limit is -infinity. Does all this make sense ?

OpenStudy (anonymous):

Sorry for the late answer I was working haha. and Yes it does make more sense, as x approaches 2, the number will be increasingly bigger but because the limit would be over 0 it's an asymptote if I understood well enough :)